OpenStudy (bipolarowls):

Karen is writing a word problem that can be represented by the equation 3x + 2 = 20. A box weighs 3 pounds and several 2 pound books are put in the box. The total weight of the box and books is 20 pounds. How many books are in the box? Does her word problem fit the equation? Justify your response.

OpenStudy (bipolarowls):

@santana

OpenStudy (christopher1448):

its my names jeff!

OpenStudy (trinity_weston):

x=6

OpenStudy (bipolarowls):

I will report.

OpenStudy (bipolarowls):

Please explain ;)

OpenStudy (christopher1448):

.......

OpenStudy (christopher1448):

MY NAME IS JEFF

OpenStudy (bipolarowls):

......

OpenStudy (christopher1448):

...............

OpenStudy (lxdrpr):

wouldnt the variable be on two because you dont know how many books are in the box but the box weight is known?

OpenStudy (bipolarowls):

idk that is the question D:

OpenStudy (mhchen):

No it does NOT fit. "The box weights 3 pounds" so that would be the CONSTANT. Each additional book is 2 pounds, so 'x' represents the changable value of the number of books.

OpenStudy (lxdrpr):

yay i was right

OpenStudy (bipolarowls):

Okay x= number of books

OpenStudy (mhchen):

Basically, if you have a word problem, whatever one that 'changes' over an interval or time would be the variable

OpenStudy (bipolarowls):

Omq Thanks!!!

OpenStudy (trinity_weston):

ok well if you have 3x+2=20 the 3 times the amount of x which is 6 would equal 18 +2=20

OpenStudy (mhchen):

The question is whether the equation fits the word problem. It should be 3+2x=20, not 2 + 3x=20

OpenStudy (bipolarowls):

Thanks so much!!

OpenStudy (lxdrpr):

your answer doesnt match the question lol @Trinity_Weston

OpenStudy (bipolarowls):

?